OpenStudy (anonymous):

find the slope of the curve 4x^2y+16 at the point (2,1)

OpenStudy (anonymous):

Find its derivative, 4[2x^2y(y/x + log(x)dy/dx)] (8yx^2y/x) + (8x^2y.log(x) dy/dx) = dy/dx (8yx^2y/x) = dy/dx(1 - 8x^2y.log(x) ) dy/dx = (8yx^2y/x) /(1 - 8x^2y.log(x)) So putting the values of x and y (2,1) dy/dx = [(8)(1)(2^2)/2]/[1 - (8)(2^2)(log(2))] = 16 / (1-32 log 2) =16/-21.1 =-0.755